# Geodesic Calculus on Latent Spaces

### v1

@@ Abstract.
Previous work has focused on learning underlying structures, e.g., manifold representations (Arvanitidis et al., 2018) or Riemannian metrics (Gruffaz & Sassen, 2025). However, practical computation of geodesic interpolation between given endpoints or geodesic extrapolation remains challenging. We address this by making the latent manifold’s geometry accessible via an implicit representation based on a learned projection that minimizes a denoising objective. We further introduce a time-discrete variational geodesic calculus suitable for imperfect implicit representations, along with practical computational algorithms. Figure˜1 shows an illustrative example for both components. Building on time discretizations proven effective for Riemannian shape spaces (Rumpf & Wirth, 2015), we hope to open up new ways of using latent representations in machine learning in general and enable new reduced-order methods for shape spaces in particular.
@@
  • $\circ$
    We evaluate our approach on various autoencoders trained on synthetic and real data.
@@
Their widespread use has made the latent space of autoencoders a prime object of study, and equipping them with an appropriate notion of geometry is a major goal. Shao et al. (2018) compute interpolations and other geometric operations on the data manifold by parametrizing it via the decoder, which is equivalent to pulling back the Euclidean metric from the data space to the latent space. However, they disregard any non-Euclidean geometric structure of the latent space. Chen et al. (2018) pursue a similar idea for VAEs, where they additionally modify the pulled-back metric to generate high cost away from the data manifold. This underlying idea was concurrently pursued by Arvanitidis et al. (2018) based on previous work by Tosi et al. (2014) on Gaussian process latent variable models. They further extended this idea to pulling back non-Euclidean metrics (Arvanitidis et al., 2021), to pulling back the Fisher–Rao metric on densities (Arvanitidis et al., 2022; Lobashev et al., 2025), and to using Finsler metrics on latent spaces (Pouplin et al., 2023). In contrast to these works, we propose to encode the latent manifold as an implicit submanifold and derive a discrete geodesic calculus from this description. Sun et al. (2025) also learn an implicit representation of the latent manifold, however, they use it to modify the metric on the latent space and pull this modified metric back to the data space via the encoder. We will perform all optimizations directly on the latent manifold to maintain the advantages of its low dimensionality.
@@
One of the most fundamental tasks in Riemannian geometry is the computation of geodesic paths, either solving the system of geodesic differential equations via numerical integration techniques for given initial data, as in (Beg et al., 2005), or minimizing the so-called path energy over paths with prescribed endpoints. The latter variant is based on a suitable discretization of the path energy depending on the type of Riemannian space. In section˜3 we will follow this time discretization paradigm. By (Rumpf & Wirth, 2015) it was developed into a comprehensive time-discrete geodesic calculus on shape spaces (including a corresponding convergence analysis for vanishing time steps), and it was applied in different contexts such as curves (Bauer et al., 2017), discrete surfaces (Heeren et al., 2014), or images (Berkels et al., 2015).
@@
In this paper, we consider an autoencoder for data ${\mathcal{X}}\subset{\mathcal{M}}\subset\mathbb{R}^{n}$ on a hidden manifold $\mathcal{M}$} consisting of an encoder $\phi\colon\mathbb{R}^{n}\to\mathbb{R}^{l}$ and decoder $\psi\colon\mathbb{R}^{l}\to\mathbb{R}^{n}$. We will describe the corresponding latent manifold ${\mathcal{Z}}\coloneqq\phi({\mathcal{M}})$ as an implicit submanifold with a Riemannian metric and develop suitable numerical schemes for geodesic interpolation and extrapolation. Let us shortly recap the basics of the required Riemannian geometry: We consider the manifold ${\mathcal{Z}}\subset\mathbb{R}^{l}$ as an implicit, $m$-dimensional submanifold, i.e. we have ${\mathcal{Z}}\coloneqq\left\{{z}\in\mathbb{R}^{l}\mid\zeta({z})=0\right\}$ for a smooth map $\zeta$. In our case, $\zeta\colon\mathbb{R}^{l}\to\mathbb{R}^{l};\;\zeta({z})\coloneqq{z}-\Pi({z})$, where $\Pi\colon\mathbb{R}^{l}\to\mathbb{R}^{l}$ is the projection from the latent space $\mathbb{R}^{l}$ to the nearest point on the latent manifold $\mathcal{Z}$}. The tangent space $T_{z}{\mathcal{Z}}$ at a point ${z}\in{\mathcal{Z}}$ is the vector space of all velocities of paths passing through z{z}. For implicit submanifolds, $T_{z}{\mathcal{Z}}=\ker D\zeta({z})$. A Riemannian metric is an inner product $g_{z}(\cdot,\cdot)$ on $T_{z}{\mathcal{Z}}$ smoothly depending on z{z}. This inner product allows to measure lengths of tangent vectors and angles between them. The simplest choice would be the Euclidean inner product inherited from $\mathbb{R}^{l}$, but it may be useful to apply other inner products that better represent the geometric structure of the original data.
The length of a path $\mathbf{{z}}\colon[0,1]\to{\mathcal{Z}}$ with velocity $\dot{\mathbf{{z}}}$ is defined as ${\mathcal{L}}(\mathbf{{z}})=\int_{0}^{1}\sqrt{g_{\mathbf{{z}}(t)}(\dot{\mathbf{{z}}}(t),\dot{\mathbf{{z}}}(t))}\,\,\mathrm{d}t$, and the Riemannian distance $\mathrm{dist}({z}_{0},{z}_{1})$ between two points ${z}_{0},{z}_{1}\in{\mathcal{Z}}$ is the infimum over all paths with endpoints $\mathbf{{z}}(0)={z}_{0}$, $\mathbf{{z}}(1)={z}_{1}$. A minimizing path is called a geodesic. A torus $\mathcal{Z}$} is shown as a toy example for $l=3$ in fig.˜2. A geodesic connecting ${z}_{0}$ and ${z}_{1}$ can equivalently be found by minimizing the path energy
@@
of a discrete path $\mathbf{{z}}=({z}_{0},\ldots,{z}_{K})$ as a discrete counterpart of (1). Consequently, a discrete geodesic for given endpoints ${z}_{0},{z}_{K}\in{\mathcal{Z}}$ is a minimizer of (2) subject to the constraint $\zeta({z}_{k})=0$ for $k=0,\ldots,K$. Rumpf & Wirth (2015, Corollary 4.10) have shown that if $\mathcal{Z}$} is smooth, ${z}\mapsto g_{z}$ is Lipschitz continuous, and $\mathcal{W}({z}_{0},{z}_{1})$ approximates $\mathrm{dist}({z}_{0},{z}_{1})^{2}$ up to an error $O(\,\mathrm{dist}({z}_{0},{z}_{1})^{-3})$, then the piecewise affine interpolations of minimizers of the discrete path energies ${\mathcal{E}}^{K}$ converge uniformly to minimizers of the continuous path energy $\mathcal{E}$}. One can interpret $\mathcal{W}({z}_{k-1},{z}_{k})$ physically as the energy of a spring connecting ${z}_{k-1}$ and ${z}_{k}$. Then, ${\mathcal{E}}^{K}({z}_{0},\ldots,{z}_{K})$ is the total elastic energy of a chain of $K$ springs, which we relax under the constraint that all nodes lie on $\mathcal{Z}$}. Depending on the application and the underlying configuration of the data manifold, one can distinguish different Riemannian metrics on $\mathcal{Z}$} and corresponding functionals $\mathcal{W}$, for instance,
  • $\circ$
@@
    (3) $\mathcal{W}_{\mathrm{E}}({z}_{0},{z}_{1})\coloneqq|{z}_{1}-{z}_{0}|^{2},$
  • $\circ$
    the pullback metric $g_{z}(v,w)=g^{\mathcal{M}}_{\psi({z})}(D\psi({z})v,D\psi({z})w)$ pulling back the metric
    $g^{\mathcal{M}}_{x}:T_{x}{\mathcal{M}}\times T_{x}{\mathcal{M}}\to\mathbb{R}$ of $\mathcal{M}$} results in
    (4) $),\displaystyle\mathcal{W}_{\mathcal{M}}({z},\tilde{z})=\,\mathrm{dist}^{2}_{{\mathcal{M}}}(\psi({z}),\psi(\tilde{z}$)),
  • $\circ$
@@
for $k=1,\ldots,K-1$ and $i=1,\ldots,l$.
We propose to use an augmented Lagrangian method to compute solutions of the constrained optimization problem. In detail, for the augmented Lagrangian
(8) ${\mathbf{L}}^{a}(\mathbf{{z}}_{j},\Lambda_{j},\mu_{j})={\mathbf{L}}(\mathbf{{z}}_{j},\Lambda_{j})+\tfrac{\mu_{j}}{2}\,|\zeta(\mathbf{{z}}_{j})|^{2}={\mathcal{E}}(\mathbf{{z}}_{j})-\Lambda_{j}:\zeta(\mathbf{{z}}_{j})+\tfrac{\mu_{j}}{2}\,|\zeta(\mathbf{{z}}_{j})|^{2}$
@@
The augmented Lagrangian approach is not harmed by the fact that the Lagrange multiplier in (6)-(7) is underdetermined (and thus nonunique) due to $\text{rank\;}D\zeta({z})=l-m$. Moreover, it is also applicable with inexact constraints (Frick et al., 2011; Jin, 2017), which is important since in practice we replace $\zeta$ with $\zeta_{\sigma}=\mathrm{id}-\Pi_{\sigma}$ for a learned approximate projection $\Pi_{\sigma}$ (cf. section˜4). The augmented Lagrangian method allows this approximation as long as $\zeta_{\sigma}$ points approximately in normal direction to $\mathcal{Z}$}. Furthermore, the penalty term enforces small values of $\zeta_{\sigma}({z}_{k})={z}_{k}-\Pi_{\sigma}({z}_{k})$ even though $\zeta_{\sigma}$ is not expected to vanish. Overall, we observe that the augmented Lagrangian method works for an inexact implicit function $\zeta_{\sigma}$ as long as $(\zeta_{\sigma},D\zeta_{\sigma})$ approximate $(\zeta,D\zeta)$ sufficiently well (see, e.g., fig.˜2). We give details on the implementation and the parameters for the augmented Lagrangian approach in section˜A.2.
@@
To derive a discrete counterpart of geodesic extrapolation via the exponential map, we interpret the first saddle point condition
$0=K(\partial_{{z}_{k}}\mathcal{W}({z}_{k-1},{z}_{k})+\partial_{{z}_{k}}\mathcal{W}({z}_{k},{z}_{k+1}))-\nabla\zeta({z}_{k})\lambda_{k}$
@@
(11) $\mathcal{F}({z}_{k+1},\lambda_{k})$ $\coloneqq|K(\partial_{{z}_{k}}\mathcal{W}({z}_{k-1},{z}_{k})+\partial_{{z}_{k}}\mathcal{W}({z}_{k},{z}_{k+1}))-\nabla\zeta({z}_{k})\lambda_{k}|^{2}+\tfrac{\mu}{2}|\zeta({z}_{k+1})|^{2},$
where $\mu>0$ is some penalty parameter. As in the context of the discrete geodesic interpolation, this minimization remains reasonable if $\zeta$ is replaced by an approximation $\zeta_{\sigma}$. Now, given the first two points ${z}_{0}$ and ${z}_{1}$ and thus a corresponding discrete initial velocity $v_{0}=K({z}_{1}-{z}_{0})$, we iteratively minimize $\mathcal{F}({z}_{k+1},\lambda_{k})$ for $k=1,\ldots,K-1$ in both ${z}_{k+1}$ and $\lambda_{k}$ with a BFGS method. We define by $\mathrm{Exp}^{K}_{{z}_{0}}(v_{0})\coloneqq{z}_{K}$ the discrete exponential map as the final point of the extrapolated discrete geodesic $({z}_{0},\ldots,{z}_{K})$ (cf. fig.˜3). Following Rumpf & Wirth (2015, Theorem 5.1), the discrete exponential map converges to the continuous one as $K\to\infty$.
Figure 3. Comparison between computed exponentials with learned implicit manifold representation $\zeta_{\sigma}$ (green points) and ground truth representation $\zeta$ (black line). As in any dynamical system, slight numerical inaccuracies lead to an exponentially growing divergence (which is known to be more pronounced in regions of negative curvature as in the right-most example).
@@
Let us suppose a density measure $\mathrm{d}z$ is given on the latent manifold $\mathcal{Z}$} reflecting the sampling on the data manifold. In applications, this is typically the empirical measure of the data samples or rather of their images under encoder $\phi$. The projection as a neural network is learned based on this possibly noisy point cloud representing the latent manifold.
Following the approach to learn a projection from denoising autoencoders proposed by Alain & Bengio (2014), we define the projection $\Pi_{\sigma}$ as the minimizer of the loss functional
@@
The submanifold is designed to approximate a dataset of shapes, such as different poses of a humanoid model. Because many datasets provide only a limited number of examples, we first apply a classical Riemannian construction to obtain the submanifold and its parametrization, which is computationally demanding. We then use this parametrization to create a denser training set for our autoencoder. We discard any pairs where at least one shape exhibits self-intersections to preserve physical plausibility. See section˜A.4, for details on the data generation and the autoencoder training. In this way, the autoencoder learns a low-dimensional latent manifold approximating the shape space submanifold, enabling us to perform discrete geodesic operations in reduced dimensions. Figure˜1 illustrates the overall procedure of our approach for an example of a two-dimensional submanifold of the manifold of discrete shells.
@@
As described in section˜4, we use the trained projection operator $\Pi_{\sigma}$ to construct an approximate implicit representation of the latent manifold. For the geodesic calculus, we use the Euclidean metric, as this agrees with the shell distance due to the autoencoder’s isometry. In fig.˜4, we show a result of applying this approach to the SCAPE dataset (Anguelov et al., 2005) of human character poses and compare linear interpolation in latent space with geodesics on the latent manifold computed using our approach. Our geodesics avoid self-intersections more effectively than linear interpolation, thanks to the rejection of intersecting shapes during training. Since our learned projection $\Pi_{\sigma}$ is only accurate up to the filter width $\sigma$, slight self-intersections may remain. We show additional examples in section˜A.4.
@@
We consider a latent manifold resulting from training a spherical variational autoencoder (SVAE) Davidson et al. (2018) on motion capture data from the CMU Graphics Lab . A suitable representation for the data was described by Tournier et al. (2009): A pose is defined as an element of $\mathrm{SO}(3)^{m}$, where $m=30$ is the number of joints in the skeleton. The vector of rotations specifies the rotations of the joints. Given the connectivity and lengths of the skeletal segments, the full pose can be reconstructed. Hence, our input data $\mathcal{X}$} lies on a hidden manifold ${\mathcal{M}}\subset\mathrm{SO}(3)^{m}$. Arvanitidis et al. (2022) used an SVAE autoencoder to take the hyperspherical nature of this data into account. Unlike standard VAEs, which assume a Gaussian prior in the latent space, the SVAE uses von Mises–Fisher (vMF) distributions for regularization, which indeed enforces a hyperspherical latent geometry. Details on the employed data and training are provided in section˜A.5. We use $l=10$ latent dimensions. In the case of (S)VAEs, the encoder and decoder maps are not deterministic but parameterize distributions. We sample the latent manifold $\mathcal{Z}$} by sampling from the encoder distribution and obtain decoded points by sampling from the decoder. For simplicity, when learning $\Pi_{\sigma}$ and calculating discrete geodesics on $\mathcal{Z}$}, we ignore the nondeterministic nature of the encoder and decoder and, by a slight abuse of notation, denote by $\psi({z})\in\mathrm{SO}(3)^{m}$ a sample of the decoding of z{z}. Another approach would be to follow Arvanitidis et al. (2022) and incorporate the Kullback–Leibler divergence, see ˜14 in section˜A.1.
In figs.˜5 and 6 (left), we show a projection of the sampled latent manifold (from which we learn $\Pi_{\sigma}$) onto the three most relevant dimensions obtained from a PCA.
@@
(13) $\mathcal{W}_{{\mathcal{M}}}({z},\tilde{z})=\,\mathrm{dist}^{2}_{\mathrm{SO}(3)}(\psi({z}),\psi(\tilde{z}))=\sum_{i=1}^{m}\left|\arccos\!\left(\psi({z})_{i}\cdot\psi(\tilde{z})_{i}\right)\right|^{2}.$
In fig.˜5 (right), we compare geodesic interpolation based on $\mathcal{W}_{{\mathcal{M}}}$ (green) to geodesic interpolation with the Euclidean metric $\mathcal{W}_{\text{E}}({z},\tilde{z})=|{z}-\tilde{z}|^{2}$ (yellow) as well as linear interpolation in latent space (red). The results show that geodesic interpolation with the pullback metric yields realistic decoded paths, whereas linear interpolation in latent space leads to poses not lying on the data manifold $\mathcal{M}$}, as indicated by unnatural contractions of shoulders and hips and by self-intersecting limbs. Figure˜6 shows a pose extrapolation using the exponential map on the latent manifold. Different from linear extrapolation, the extrapolated path follows the geometry of the latent manifold and, after decoding, leads to realistic poses.

### v2

# Geodesic Calculus on Implicitly Defined Latent Manifolds

@@ Abstract.
Previous work has focused on learning underlying structures, e.g., manifold representations (Arvanitidis et al., 2018) or Riemannian metrics (Gruffaz & Sassen, 2025). The computation of geodesics on latent manifolds is often restricted to specific autoencoders (Arvanitidis et al., 2022) and specific choices of the metric. Practical computation of geodesic interpolation between given endpoints or geodesic extrapolation for some initial position and direction that can be seamlessly integrated in an existing autoencoder pipeline remains challenging. There are two essential properties to be achieved. First, the computed paths must adhere to the latent sample distribution. Second, for non-isometric embeddings a problem-dependent metric must be additionally incorporated. To address this, we make the latent manifold’s geometry accessible via an implicit representation based on a learned projection that minimizes a denoising objective. Furthermore, we introduce a time-discrete variational geodesic calculus which allows for different metrics and is suitable for imperfect implicit representations. This tool can be used in diverse applications and is already known to approximate ground truth geodesics well in the case of perfect implicit representations. Figure˜1 shows an illustrative example for both desired properties. In addition the toy example in fig.˜2 motivates the need of an actual representation of the latent manifold geometry for geodesic interpolation purposes. Building on a suitable time discretization which proved effective for Riemannian shape spaces (Rumpf & Wirth, 2015), we hope to open up new ways of using latent representations in machine learning in general and enable new reduced-order methods for shape spaces in particular.
@@
  • We evaluate our approach on various autoencoders trained on synthetic and on real data.
@@
Their widespread use has made the latent space of autoencoders a prime object of study, and equipping them with an appropriate notion of geometry is a major goal. Shao et al. (2018) compute interpolations and other geometric operations on the data manifold by parametrizing it via the decoder. However, they ignore that the actual latent manifold may have nonzero codimension in latent space. Chen et al. (2018) pursue a similar idea for VAEs, where they additionally modify the pulled-back metric to generate high cost away from the data manifold. This underlying idea was concurrently pursued by Arvanitidis et al. (2018) based on previous work by Tosi et al. (2014) on Gaussian process latent variable models. They further extended this idea to pulling back non-Euclidean metrics (Arvanitidis et al., 2021), to pulling back the Fisher–Rao metric on densities (Arvanitidis et al., 2022; Lobashev et al., 2025a), and to using Finsler metrics on latent spaces (Pouplin et al., 2023). In contrast to these works, we propose to encode the latent manifold as an implicit submanifold and derive a discrete geodesic calculus from this description. This allows both to compute paths that follow the latent manifold, i.e. the data support, and also to incorporate application-dependent metrics. Sun et al. (2025) also learn an implicit representation of the latent manifold, however, they use it to modify the metric on the latent space and pull this modified metric back to the data space via the encoder. We will perform all optimizations directly on the latent manifold to maintain the advantages of its low dimensionality. In section˜C.3 and section˜C.4 we further discuss and compare with different methods that compute paths following the data density.
@@
One of the most fundamental tasks in Riemannian geometry is the computation of geodesic paths, either solving the system of geodesic differential equations via numerical integration techniques for given initial data, as in (Beg et al., 2005), or minimizing the so-called path energy over paths with prescribed endpoints. The latter variant is based on a suitable discretization of the path energy depending on the type of Riemannian space. In section˜3 we will follow this time discretization paradigm. By Rumpf & Wirth (2015) it was developed into a comprehensive time-discrete geodesic calculus on shape spaces (including a corresponding convergence analysis for vanishing time steps), and it was applied in different contexts such as curves (Bauer et al., 2017), discrete surfaces (Heeren et al., 2014), or images (Berkels et al., 2015). In section˜C.1 and section˜C.2 we more extensively compare different manifold representations and numerical discretizations.
@@
Figure 2. Standard autoencoder trained on images of a rotated and scaled letter ‘A’ with two-dimensional latent space. Comparison between geodesic interpolation (green) using our method and linear interpolation (red) showing below every third decoded image.
@@
In this paper, we consider an autoencoder for data ${\mathcal{X}}\subset{\mathcal{M}}\subset\mathbb{R}^{n}$ on a hidden manifold $\mathcal{M}$} consisting of an encoder $\phi\colon\mathbb{R}^{n}\to\mathbb{R}^{l}$ and decoder $\psi\colon\mathbb{R}^{l}\to\mathbb{R}^{n}$. We will describe the corresponding latent manifold ${\mathcal{Z}}\coloneqq\phi({\mathcal{M}})$ as an implicit submanifold with a Riemannian metric. This enables us to develop suitable numerical schemes for geodesic interpolation and extrapolation without requiring an explicit parametrization. Let us shortly recap the basics of the required Riemannian geometry: We consider the manifold ${\mathcal{Z}}\subset\mathbb{R}^{l}$ as an implicit, $m$-dimensional submanifold, i.e. we have ${\mathcal{Z}}\coloneqq\left\{{z}\in\mathbb{R}^{l}\mid\zeta({z})=0\right\}$ for a smooth map $\zeta$. In our case,
(1) $\zeta\colon\mathbb{R}^{l}\to\mathbb{R}^{l};\;\zeta({z})\coloneqq{z}-\Pi({z}),$
where $\Pi\colon\mathbb{R}^{l}\to\mathbb{R}^{l}$ is the projection from the latent space $\mathbb{R}^{l}$ to the nearest point on the latent manifold $\mathcal{Z}$}. The tangent space $T_{z}{\mathcal{Z}}$ at a point ${z}\in{\mathcal{Z}}$ is the vector space of all velocities of paths passing through z{z}. For implicit submanifolds, $T_{z}{\mathcal{Z}}=\ker D\zeta({z})$. A Riemannian metric is an inner product $g_{z}(\cdot,\cdot)$ on $T_{z}{\mathcal{Z}}$ smoothly depending on z{z}. This inner product allows to measure lengths of tangent vectors and angles between them. The simplest choice would be the Euclidean inner product inherited from $\mathbb{R}^{l}$, but it may be useful to apply other inner products that better represent the geometric structure of the original data.
The length of a path $\mathbf{{z}}\colon[0,1]\to{\mathcal{Z}}$ with velocity $\dot{\mathbf{{z}}}$ is defined as ${\mathcal{L}}(\mathbf{{z}})=\int_{0}^{1}\sqrt{g_{\mathbf{{z}}(t)}(\dot{\mathbf{{z}}}(t),\dot{\mathbf{{z}}}(t))}\,\,\mathrm{d}t$, and the Riemannian distance $\mathrm{dist}({z}_{0},{z}_{1})$ between two points ${z}_{0},{z}_{1}\in{\mathcal{Z}}$ is the infimum over all paths with endpoints $\mathbf{{z}}(0)={z}_{0}$, $\mathbf{{z}}(1)={z}_{1}$. A minimizing path is called a geodesic. A torus $\mathcal{Z}$} is shown as a toy example for $l=3$ and $m=2$ in fig.˜3. A geodesic connecting ${z}_{0}$ and ${z}_{1}$ can equivalently be found by minimizing the path energy
@@
of a discrete path $\mathbf{{z}}=({z}_{0},\ldots,{z}_{K})$ as a discrete counterpart of (2). Consequently, a discrete geodesic for given endpoints ${z}_{0},{z}_{K}\in{\mathcal{Z}}$ is a minimizer of (3) subject to the constraint $\zeta({z}_{k})=0$ for $k=0,\ldots,K$. Rumpf & Wirth (2015, Corollary 4.10) have shown that if $\mathcal{Z}$} is smooth, ${z}\mapsto g_{z}$ is Lipschitz continuous, and $\mathcal{W}({z}_{0},{z}_{1})$ approximates $\mathrm{dist}({z}_{0},{z}_{1})^{2}$ up to an error $O(\,\mathrm{dist}({z}_{0},{z}_{1})^{3})$, then the piecewise affine interpolations of minimizers of the discrete path energies ${\mathcal{E}}^{K}$ converge uniformly to minimizers of the continuous path energy $\mathcal{E}$}. One can interpret $\mathcal{W}({z}_{k-1},{z}_{k})$ physically as the energy of a spring connecting ${z}_{k-1}$ and ${z}_{k}$. Then, ${\mathcal{E}}^{K}({z}_{0},\ldots,{z}_{K})$ is the total elastic energy of a chain of $K$ springs, which we relax under the constraint that all nodes lie on $\mathcal{Z}$}. Depending on the application and the underlying configuration of the data manifold, one can distinguish different Riemannian metrics on $\mathcal{Z}$} and corresponding functionals $\mathcal{W}$:
@@
    (4) $\mathcal{W}_{\mathrm{E}}({z},\tilde{z})=|\tilde{z}-{z}|^{2}.$
  • The pullback metric (in the case of an explicitly known metric on $\mathcal{M}$})
    $g_{z}(v,w)=g^{\mathcal{M}}_{\psi({z})}(D\psi({z})v,D\psi({z})w)$ with $g^{\mathcal{M}}_{x}:T_{x}{\mathcal{M}}\times T_{x}{\mathcal{M}}\to\mathbb{R}$ is reflected by
    (5) $).\displaystyle\mathcal{W}_{\mathcal{M}}({z},\tilde{z})=\,\mathrm{dist}^{2}_{{\mathcal{M}}}(\psi({z}),\psi(\tilde{z}$)).
@@
for $k=1,\ldots,K-1$ and $i=1,\ldots,l$. We propose to use an augmented Lagrangian method to compute solutions of the constrained optimization problem. In detail, for the augmented Lagrangian
(9) ${\mathbf{L}}^{a}(\mathbf{{z}}_{j},\Lambda_{j},\mu_{j})={\mathbf{L}}(\mathbf{{z}}_{j},\Lambda_{j})+\tfrac{\mu_{j}}{2}\,|\zeta(\mathbf{{z}}_{j})$ $={\mathcal{E}}^{K}(\mathbf{{z}}_{j})-\Lambda_{j}:\zeta(\mathbf{{z}}_{j})+\tfrac{\mu_{j}}{2}\,|\zeta(\mathbf{{z}}_{j})|^{2}$
@@
The augmented Lagrangian approach is not harmed by the fact that the Lagrange multiplier in (7)-(8) is underdetermined (and thus nonunique) due to $\text{rank\;}D\zeta({z})=l-m$. Moreover, it is also used with inexact constraints (Frick et al., 2011; Jin, 2017), which is important since in practice we replace $\zeta$ with $\zeta_{\sigma}=\mathrm{id}-\Pi_{\sigma}$ for a learned approximate projection $\Pi_{\sigma}$ (cf. section˜4). The augmented Lagrangian method allows this approximation as long as $\zeta_{\sigma}$ points approximately in normal direction to $\mathcal{Z}$}. Furthermore, the penalty term enforces small values of $\zeta_{\sigma}({z}_{k})={z}_{k}-\Pi_{\sigma}({z}_{k})$ even though $\zeta_{\sigma}$ is not expected to vanish. Overall, we observe that the augmented Lagrangian method works for an inexact implicit function $\zeta_{\sigma}$ as long as $(\zeta_{\sigma},D\zeta_{\sigma})$ approximate $(\zeta,D\zeta)$ sufficiently well (see, e.g., fig.˜3). We give details on the implementation and the parameters for the augmented Lagrangian approach in section˜A.2.
@@
To derive a discrete counterpart of geodesic extrapolation via the exponential map, we aim at a numerical scheme, which reproduces discrete variational geodesics minimizing ˜3. To this end, we interpret the first saddle point condition
$\partial_{{z}_{k}}\widehat{\mathbf{L}}(({z}_{k-1},{z}_{k},{z}_{k+1}),\lambda_{k})\coloneqq\partial_{{z}_{k}}{\mathbf{L}}(\mathbf{{z}},\Lambda)=0$
@@
(12) $\mathcal{F}({z}_{k+1},\lambda_{k})\coloneqq$ $\tfrac{\mu}{2}|\zeta({z}_{k+1})|^{2}+|\partial_{{z}_{k}}\widehat{\mathbf{L}}(({z}_{k-1},{z}_{k},{z}_{k+1}),\lambda_{k})|^{2},$
where $\mu>0$ is some penalty parameter. As in the context of the discrete geodesic interpolation, this minimization remains reasonable if $\zeta$ is replaced by an approximation $\zeta_{\sigma}$. Now, given the first two points ${z}_{0}$ and ${z}_{1}$ and thus a corresponding discrete initial velocity $v_{0}=K({z}_{1}-{z}_{0})$, we iteratively minimize $\mathcal{F}({z}_{k+1},\lambda_{k})$ for $k=1,\ldots,K-1$ in both ${z}_{k+1}$ and $\lambda_{k}$ with a BFGS method. We define by $\mathrm{Exp}^{K}_{{z}_{0}}(v_{0})\coloneqq{z}_{K}$ the discrete exponential map as the final point of the extrapolated discrete geodesic $({z}_{0},\ldots,{z}_{K})$ (cf. fig.˜4). Following Rumpf & Wirth (2015, Theorem 5.10), the discrete exponential map converges to the continuous one as $K\to\infty$.
Functional ˜12 involves first derivatives of the distance approximation $\mathcal{W}$ and the implicit representation $\zeta$. However, these derivatives are not taken with respect to the optimization variable ${z}_{k+1}$. As a consequence, when computing the gradient of (12), the network representing $\zeta$ does not need to be differentiated twice, and only mixed second derivatives of $\mathcal{W}$ are required. Even when $\mathcal{W}$ is defined via a pullback metric, these derivatives can be computed without second-order differentiation of the decoder network.
Figure 4. Comparison between computed exponentials with learned implicit manifold representation $\zeta_{\sigma}$ (green points) and ground truth representation $\zeta$ (black line). As in any dynamical system, slight numerical inaccuracies lead to an exponentially growing divergence (which is known to be more pronounced in regions of negative curvature as in the right-most example). See also the experiment in fig.˜15.
@@
Let us suppose a density measure $\mathrm{d}z$ is given on the latent manifold $\mathcal{Z}$} reflecting the sampling on the data manifold. In applications, this is typically the empirical measure of the data samples or rather of their images under encoder $\phi$. The projection as a neural network is learned based on this possibly noisy point cloud representing the latent manifold. A natural approach is to learn it via the variational approach proposed by Alain & Bengio (2014) in the context of denoising autoencoders: We define the projection $\Pi_{\sigma}$ as the minimizer of the loss functional
@@
The submanifold is designed to approximate a dataset of shapes, such as different poses of a humanoid model. Because many datasets provide only a limited number of examples, we first apply a classical Riemannian construction to obtain the submanifold and its parametrization, which is computationally demanding. We then use this parametrization to create a denser training set for our autoencoder. We discard any pairs where at least one shape exhibits self-intersections to preserve physical plausibility. See section˜B.1, for details on the data generation and the autoencoder training. In this way, the autoencoder learns a low-dimensional latent manifold approximating the shape space submanifold, enabling us to perform discrete geodesic operations in reduced dimensions. Figure˜1 illustrates the overall procedure of our approach for an example of a two-dimensional submanifold of the manifold of discrete shells.
@@
As described in section˜4, we use the trained projection operator $\Pi_{\sigma}$ to construct an approximate implicit representation of the latent manifold. For the geodesic calculus, we use the Euclidean metric, as this agrees with the shell distance due to the autoencoder’s isometry. In fig.˜5, we show a result of applying this approach to the SCAPE dataset (Anguelov et al., 2005) of human character poses and compare linear interpolation in latent space with geodesics on the latent manifold computed using our approach. Our geodesics avoid self-intersections more effectively than linear interpolation, thanks to the rejection of intersecting shapes during training. Since our learned projection $\Pi_{\sigma}$ is only accurate up to the filter width $\sigma$, slight self-intersections may remain. We show additional examples in section˜B.1.
@@
We consider a latent manifold resulting from training a spherical variational autoencoder (SVAE) Davidson et al. (2018) on motion capture data from the CMU Graphics Lab . A suitable representation for the data was described by Tournier et al. (2009): A pose can be represented as an element of $\mathrm{SO}(3)^{m}$, where $m=30$ is the number of joints in the skeleton. Here, we identify rotation matrices with corresponding rotation vectors. The vector of rotations specifies the rotations of the joints. Given the connectivity and lengths of the skeletal segments, the full pose can be reconstructed. Hence, our input data $\mathcal{X}$} lies on a hidden manifold ${\mathcal{M}}\subset\mathrm{SO}(3)^{m}$. Arvanitidis et al. (2022) used an SVAE autoencoder to take the hyperspherical nature of this data into account. Unlike standard VAEs, which assume a Gaussian prior in the latent space, the SVAE uses von Mises–Fisher (vMF) distributions for regularization, which indeed enforces a hyperspherical latent geometry. Details on the employed data and training are provided in section˜B.2. We use $l=10$ latent dimensions. In the case of (S)VAEs, the encoder and decoder maps are not deterministic but parameterize distributions. We sample the latent manifold $\mathcal{Z}$} by sampling from the encoder distribution and obtain decoded points by sampling from the decoder. For simplicity, when learning $\Pi_{\sigma}$ and calculating discrete geodesics on $\mathcal{Z}$}, we ignore the nondeterministic nature of the encoder and decoder and, by a slight abuse of notation, denote by $\psi({z})\in\mathrm{SO}(3)^{m}$ the mean of the distribution obtained from decoding z{z}. Another approach would be to follow Arvanitidis et al. (2022) and incorporate the Kullback–Leibler divergence, see ˜14 in section˜A.1 and section˜C.3.
In fig.˜6 (left), we show a projection of the sampled latent manifold (from which we learn $\Pi_{\sigma}$) onto the three most relevant dimensions obtained from a PCA.
@@
$\mathcal{W}_{{\mathcal{M}}}({z},\tilde{z})=\,\mathrm{dist}^{2}_{\mathrm{SO}(3)}(\psi({z}),\psi(\tilde{z}))=\sum_{i=1}^{m}\left|\arccos\!\left(\psi({z})_{i}\cdot\psi(\tilde{z})_{i}\right)\right|^{2}.$
In fig.˜6 (right), we compare geodesic interpolation based on $\mathcal{W}_{{\mathcal{M}}}$ (green) to geodesic interpolation with the Euclidean metric $\mathcal{W}_{\text{E}}({z},\tilde{z})=|{z}-\tilde{z}|^{2}$ (yellow) as well as linear interpolation in latent space (red). The results show that geodesic interpolation with the pullback metric yields realistic decoded paths, whereas linear interpolation in latent space leads to poses not lying on the data manifold $\mathcal{M}$}, as indicated by unnatural contractions of shoulders and hips and by self-intersecting limbs. Figure˜7 shows a pose extrapolation using the exponential map on the latent manifold. Different from linear extrapolation, the extrapolated path follows the geometry of the latent manifold and, after decoding, leads to realistic poses.